Palestine Journal of Mathematics Vol. 11 (Special Issue )(2022) , 1–1

© Palestine Polytechnic University-PPU 2022

TITLE

Shalosh B. EKhad and Doron Zeilberger

Communicated by Ayman Badawi

MSC 2010 Classifications: Primary 05A30

Keywords and phrases: Matrix powers, Larcombe relations, bijective proofs, weight-enumerator

Abstract

1 Introduction 2 Section 1

Lemma 2.1. The following conditions are equivalent:

(1) S is a finite semigroup with no proper subsemigroups;

(2) S is a trivial group (that is, a group with only one element). Proof. Proposition 2.2. KAKI Proof. Corollary 2.3. KAKI Proof. KAKI Remark 2.4. KAKI

References

[1] K. E. Aubert, On the ideal theory of commutative semi-groups, Math. Scand. 1, 39–54 (1953).

Author information

Shalosh B. EKhad and Doron Zeilberger, Department of Mathematics, Rutgers University (New Brunswick) Piscataway, NH 08544, U. S. A..

E-mail: DoronZeil@gmail.com